# library.bib

### v2

@@ 1.2 Overview of paper
§2 gives background information on variations of Schur–Weyl duality, including a refinement of Schur–Weyl–Jones duality and a projection formula given by the author in [Cassidy2023]. We also give a brief overview of the Weingarten calculus for the symmetric group and a number of necessary results from [MAgeeDelaSalle].
@@ Acknowledgements
We thank Michael Magee for the many conversations regarding this paper, as well as Doron Puder, Lars Louder and Ramon van Handel for their valuable input. We also thank Noam Ta Shma for his comments and feedback on our initial version of this paper.
@@ 2.1 Symmetric group
We will denote by $S_{n}$ the symmetric group on $n$ elements and view this as the group of permutations of $[n]\overset{\mathrm{def}}{=}\{1,\dots,n\}$.
@@
The equivalence classes of complex irreducible representations of $S_{n}$ are in a natural bijection with the set of Young diagrams $\mu\vdash n,$ see [VershikOkounkov] for example. We will denote the representation corresponding to $\mu$ by $V^{\mu}$ and denote its character by $\chi^{\mu}.$
Every finite–dimensional complex representation $V$ of $S_{n}$ admits a decomposition
@@
unique up to isomorphism (ie unique up to isomorphisms that preserve each component $\left(V^{\mu}\right)^{\oplus a_{\mu}}$ of the decomposition). For each $\mu\vdash n,$ the subspace $\left(V^{\mu}\right)^{\oplus a_{\mu}}$ is called the $V^{\mu}$–isotypic component of $V.$
We will denote by $d_{\mu}\overset{\mathrm{def}}{=}\chi^{\mu}(e),$ so that $d_{\mu}=\mathrm{dim}\left(V^{\mu}\right).$ Given a representation $V$ of $S_{n}$ and a permutation $\sigma,$ we may write $V(\sigma)$ or just $\sigma$ to mean the element of $\mathrm{End}\left(V\right)$ corresponding to $\sigma,$ defined by the representation.
Given $\lambda=(\lambda_{1},\dots,\lambda_{l(\lambda)}),$ , we define $V^{\lambda^{+}(n)}$ to be the irreducible representation of $S_{n}$ (when $n-|\lambda|>\lambda_{1}$) corresponding to the YD
@@ 2.2 Partition algebra
The partition algebra appeared in the work of Martin [Martin1994], Martin and Saleur [MartinSaleur1992] and Jones [Jones] as a generalization of the Temperley–Lieb algebra and the Potts model in statistical mechanics. It is closely connected to the symmetric group via Schur–Weyl–Jones duality, which was first detailed explicitly by Jones in [Jones].
@@ 2.2.1 Set partitions
We denote by $\mathrm{Part}\left([m]\right)$ the set of set partitions of $[m]$. One may define a natural partial ordering on $\mathrm{Part}\left([m]\right)$ by refinement: $\pi_{1}\leq\pi_{2}$ if every block of $\pi_{1}$ is contained in a block of $\pi_{2}.$ We say that $\pi_{1}$ is finer than $\pi_{2}$ or that $\pi_{2}$ is coarser than $\pi_{1}.$
From this partial ordering one may derive the Möbius function $\mu$ on $\mathrm{Part}\left([m]\right)$. If $\pi_{2}=\Big{\{}\mathcal{P}_{1},\dots,\mathcal{P}_{|\pi_{2}|}\Big{\}}$ and $\pi_{1}$ is a refinement of $\pi_{2}$ in which each block $\mathcal{P}_{i}$ is split in to $b_{i}$ subsets, then
@@
as is described in the foundational paper by Rota [Rota].
@@ 2.2.2 Partition algebra
To each $\pi\in\mathrm{Part}\left([2k]\right)$ we associate a diagram consisting of two rows of $k$ vertices labeled from $1,\dots k$ on the top row and from $k+1,\dots,2k$ on the bottom row. We draw an edge between any two vertices $i$ and $j$ if and only if $i\sim j$ in $\pi.$
Any diagram consisting of two rows of $k$ vertices with all connected components complete clearly corresponds to an element $\pi\in\mathrm{Part}\left([2k]\right).$
The product $\pi_{1}\pi_{2}$ is obtained by merging the bottom row of vertices of $\pi_{1}$ with the top row of vertices of $\pi_{2}$ to obtain a new diagram $\tilde{\pi}_{3}$ with three rows of $k$ vertices; adding edges between any vertices that are in the same connected component (where there is not already an edge); removing the middle row of vertices and any adjacent edges to obtain $\pi_{3}\in\mathrm{Part}\left([2k]\right).$ Then we have
@@
Martin and Saleur [MartinSaleur1992] describe some important properties of $P_{k}(n).$
@@
The partition algebra $P_{k}(n)$ is semisimple for any $n\in\mathbb{C},$ unless $n$ is an integer in the range $[0,2k-1].$
@@ 2.3 Schur–Weyl–Jones duality
There is an analogue of classical Schur–Weyl duality due to Jones [Jones], which we refer to as Schur–Weyl–Jones duality.
@@
$\big{\langle}\left(e_{i_{i}}\otimes\dots\otimes e_{i_{k}}\right)\pi,\ e_{i_{k+1}}\otimes\dots\otimes e_{i_{2k}}\big{\rangle}=\begin{cases}1&\mathrm{if}\ j\sim l\mathrm{\ in}\ \pi\implies i_{j}=i_{l}\\ 0&\mathrm{otherwise.}\end{cases}$
@@
The symmetric group $S_{n}$ acts on $\left(\mathbb{C}^{n}\right)^{\otimes k}$ on the left via the diagonal permutation action, ie for $g\in S_{n}$ and basis vector $e_{i_{i}}\otimes\dots\otimes e_{i_{k}}\in\left(\mathbb{C}^{n}\right)^{\otimes k},$
@@ 2.4 Refinement
In [Cassidy2023], the author gives a refinement of Schur–Weyl–Jones duality. We give a brief overview of the main results of this paper. We will assume that $k\in\mathbb{Z}_{>0}$ and $n\geq 2k$.
@@
=$\overset{\mathrm{def}}{=}$ $e_{i_{1}}\otimes\dots\otimes e_{i_{j-1}}\otimes e_{i_{j+1}}\otimes\dots\otimes e_{i_{k}}.$
@@
the complex linear span of all $k$–tuples of distinct elements.
@@
$A_{k}(n)=D_{k}(n)\cap\bigcap_{j=1}^{k}\mathrm{Ker}\left(T_{j}\right).$
@@
$A_{k}(n)=\bigcap_{\pi\in\mathrm{Ker}\left(R\right)}\mathrm{Ker}\left(\pi\right).$
@@
This extends linearly to define an action of $\mathbb{C}\left[S_{k}\right]$ on $\left(\mathbb{C}^{n}\right)^{\otimes k}$ (and suitable subspaces).
We then view $A_{k}(n)$ as a representation of $S_{n}\times S_{k},$ where once again $g\in S_{n}$ acts via the diagonal permutation action and $\sigma\in S_{k}$ acts via $p_{\iota\left(\sigma^{-1}\right)}.$ We denote this representation
@@
The version of Schur–Weyl duality on $A_{k}(n)$ is given in the following theorem from [Cassidy2023].
@@
The main result of [Cassidy2023] is a formula for the orthogonal projection from $\left(\mathbb{C}^{n}\right)^{\otimes k}$ to subspaces isomorphic to the irreducible ‘blocks’ in the decomposition of $A_{k}(n).$ Define
@@
$\mathcal{U}_{\lambda,n}\overset{\mathrm{def}}{=}\big{\langle}\Delta\left(g,\sigma\right)\left(\xi_{\lambda}^{\mathrm{norm}}\right):\ g\in S_{n},\ \sigma\in S_{k}\big{\rangle}_{\mathbb{C}}\cong V^{\lambda^{+}(n)}\otimes V^{\lambda}.$
This leads to the main theorem and a corollary of the author in [Cassidy2023] which will be key for the combinatorial integration method used in this paper.
@@
$\mathcal{Q}_{\lambda,n}:\left(\mathbb{C}^{n}\right)^{\otimes k}\to\mathcal{U}_{\lambda,n}$
@@
$\sum_{\pi\in\mathrm{Part}\left([2k]\right)}c(n,k,\lambda,\pi)p_{\pi},$
@@
$c(n,k,\lambda,\pi)=d_{\lambda}d_{\lambda^{+}(n)}\frac{(-1)^{k}}{k!}\left(\sum_{\sigma,\sigma^{\prime}\in S_{k}}\sum_{\pi^{\prime}\leq\iota\left(\sigma^{-1}\sigma^{\prime}\right)\wedge\pi}\chi^{\lambda}(\sigma)\chi^{\lambda}(\sigma^{\prime})\frac{(-1)^{|\pi^{\prime}|}\mu\left(\pi^{\prime},\pi\right)}{(n)_{|\pi^{\prime}|}}\right).$ (5)
@@
$\frac{1}{d_{\lambda}}\mathrm{tr}_{\mathcal{U}_{\lambda,n}}\left(g\right)$
$=$ $\frac{1}{d_{\lambda}}\mathrm{btr}_{\left(\mathbb{C}^{n}\right)^{\otimes k}}\left(g,\mathcal{Q}_{\lambda,n}\right).$
If $\pi\leq\iota(\tau)$ for some $\tau\in S_{k},$ then the diagram corresponding to $\pi$ is obtained from that of $\iota(\tau)$ by deleting edges. Let $\mathrm{del}\left(\pi\right)$ be the number of edges deleted (ie $\mathrm{del}\left(\pi\right)=|\pi|-k$, so that if $\pi\leq\iota\left(\tau_{1}\right)$ and $\pi\leq\iota\left(\tau_{2}\right),$ then $\mathrm{del}(\pi)$ is independent of whether we count the edges deleted from $\tau_{1}$ or $\tau_{2}$).
@@
The terms in (5) with the greatest power of $n$ occur when $\pi^{\prime}=\pi.$ In this case, we have some coefficient multiplied by
$n^{k}n^{-|\pi|}=n^{k-|\pi|}=n^{-\mathrm{del}(\pi)}.$
@@ 2.5 The Weingarten calculus
An important part of the proof of Theorem 2.10 is the Weingarten calculus. Originally developed by Weingarten [Weingarten] and Xu [Xu] and then more rigorously by Collins [Collins2003] and Collins and Sniady [CollinsSniady2006], the Weingarten calculus is a method for computing integrals of matrix coefficients using Schur–Weyl duality analogues.
Below is the main formulation for the Weingarten calculus for the symmetric group, we direct the reader to the survey [CollinsWeingartenShort, Section 3] or indeed any of the aforementioned papers for an in depth discussion.
@@
For any multi–indices $I=\left(i_{1},\dots,i_{m}\right),$ $J=\left(j_{1},\dots,j_{m}\right)\in[n]^{m},$ with respect to the uniform measure, we have
@@ 2.6 $C^{*}$–algebras of free groups
We refer to [NicaSpeicher] for a thorough introduction. To the free group $F_{r}$ we associate its group algebra $\mathbb{C}\left[F_{r}\right],$ the collection of elements of the form
@@
where each $x(w)\in\mathbb{C}$ and only finitely many are non–zero.
We define a norm on $\mathbb{C}\left[F_{r}\right]$ by
@@
where $\|-\|_{\mathrm{op}}$ is the operator norm on $l^{2}\left(F_{r}\right)$. The completion of $\mathbb{C}\left[F_{r}\right]$ in this norm is the $C^{*}$–algebra of $F_{r},$ denoted $C^{*}\left(F_{r}\right)$.
The left regular representation of $F_{r}$ is denoted $\lambda$ and admits the reduced norm on $\mathbb{C}\left[F_{r}\right]$ by
@@
We define a trace $\tau$ on $\mathbb{C}\left[F_{r}\right]$ to be the map
@@ 2.7 Polynomials and random walks
We collect some results and definitions on polynomials and random walks that are required for the analysis in §3. The proofs are readily available in [MAgeeDelaSalle]. See also [CGVTvH2024] for more background.
Lemma 2.16 ([MAgeeDelaSalle, Lemma 4.2]).
@@ 2.7.1 Random walks and free groups
We call a symmetric (ie $\mu(g)=\mu\left(g^{-1}\right))$ probability measure $\mu$ on the free group $F_{r}$ reasonable if its support is finite, contains the identity element and generates $F_{r}.$
@@
The spectral radius $\rho=\rho(\mu)$ measures how fast the probability that a random walk returns to where it started decays. It is equal to the norm of $\lambda\left(\mu\right)$ on $l^{2}\left(F_{r}\right)$, where $\lambda$ is the left regular representation of $F_{r}.$
Define a proper power in the free group to be an element of the form $u^{d}$ for $u\in F_{r}$ and $d\geq 2.$
Proposition 2.18 ([MAgeeDelaSalle, Proposition 6.1]).
@@
$\mathbb{P}\left(g_{n}\ \mathrm{is\ a\ proper\ power}\right)\leq Cn^{5}\rho^{5}.$
We denote the subspace of elements of $\mathbb{C}\left[F_{r}\right]$ supported in the ball of radius $q$ by $\mathbb{C}_{\leq q}\left[F_{r}\right].$ Given a non–decreasing sequence of positive real numbers $\left(w(q)\right)_{q\in\mathbb{N}}$, we define the space
$\mathcal{S}_{w}\left(F_{r}\right)\overset{\mathrm{def}}{=}\Big{\{}x\in C^{*}\left(F_{r}\right)\ :\ x=\sum_{q}x_{q},\ x_{q}\in\mathbb{C}_{\leq q}\left[F_{r}\right],\ \sum_{q}w(q)\|x_{q}\|_{C^{*}\left(F_{r}\right)}<\infty\Big{\}}.$ (6)
If $w(q)=(1+q)^{i}$ for some $i$ then we write $\mathcal{S}_{i}\left(F_{r}\right)$ in place of $\mathcal{S}_{w}\left(F_{r}\right).$
$\mathcal{S}_{w}\left(F_{r}\right)$ is always a Banach space when equipped with the norm
$\|x\|_{w}\overset{\mathrm{def}}{=}\inf\Big{\{}\sum_{q}w(q)\|x_{q}\|_{C^{*}\left(F_{r}\right)}\ :\ x=\sum_{q}x_{q}\in\mathcal{S}_{w}\left(F_{r}\right)\Big{\}}.$

### v3

@@ 1.3 Overview of paper
§2 gives background information on variations of Schur–Weyl duality, including a refinement of Schur–Weyl–Jones duality and a projection formula given by the author in [Cas25]. We also give a brief overview of the Weingarten calculus for the symmetric group and a number of necessary results from [MdlS24].
@@ Acknowledgements
We thank Michael Magee for the many conversations regarding this paper, as well as Doron Puder, Lars Louder and Ramon van Handel for their valuable input.
@@ 2.1 Symmetric group
We will denote by $S_{n}$ the symmetric group on $n$ elements and view this as the group of permutations of $n]\overset{\mathrm{def}}{=}\{1,\dots,n\}$.
@@
The equivalence classes of finite dimensional, complex irreducible representations of $S_{n}$ are in a natural bijection with the set of Young diagrams $\mu\vdash n,$ see [VO05] for example. We will denote the representation corresponding to $\mu$ by $V^{\mu}$ and denote its character by $\chi^{\mu}.$ Every finite–dimensional complex representation $V$ of $S_{n}$ admits a decomposition
@@
unique up to isomorphism (i.e. unique up to isomorphisms that preserve each component $\left(V^{\mu}\right)^{\oplus a_{\mu}}$ of the decomposition). For each $\mu\vdash n,$ the subspace $\left(V^{\mu}\right)^{\oplus a_{\mu}}$ is called the $V^{\mu}$–isotypic component of $V.$
We will denote by $d_{\mu}\overset{\mathrm{def}}{=}\chi^{\mu}(e)=\mathrm{dim}\left(V^{\mu}\right)$. Given a representation $V$ of $S_{n}$ and a permutation $\sigma,$ we may write $V(\sigma)$ or just $\sigma$ to mean the element of $\mathrm{End}\left(V\right)$ corresponding to $\sigma,$ defined by the representation. Given $\lambda=(\lambda_{1},\dots,\lambda_{l(\lambda)}),$ , we define $V^{\lambda^{+}(n)}$ to be the irreducible representation of $S_{n}$ (when $n-|\lambda|>\lambda_{1}$) corresponding to the YD
@@ 2.2 Partition algebra
The partition algebra appeared in the work of Martin [Mar94], Martin and Saleur [MS93] and Jones [Jon94] as a generalization of the Temperley–Lieb algebra and the Potts model in statistical mechanics. It is closely connected to the symmetric group via Schur–Weyl–Jones duality, which was first detailed explicitly by Jones in [Jon94].
@@ 2.2.1 Set partitions
We denote by $\mathrm{Part}\left([m]\right)$ the set of set partitions of $[m]$. One may define a natural partial ordering on $\mathrm{Part}\left([m]\right)$ by refinement: $\pi_{1}\leq\pi_{2}$ if every block of $\pi_{1}$ is contained in a block of $\pi_{2}.$ We say that $\pi_{1}$ is finer than $\pi_{2}$ or that $\pi_{2}$ is coarser than $\pi_{1}.$ From this partial ordering one may derive the Möbius function $\mu$ on $\mathrm{Part}\left([m]\right)$. If $\pi_{2}=\Big\{\mathcal{P}_{1},\dots,\mathcal{P}_{|\pi_{2}|}\Big\}$ and $\pi_{1}$ is a refinement of $\pi_{2}$ in which each block $\mathcal{P}_{i}$ is split in to $b_{i}$ subsets, then
@@
as is described in the foundational paper by Rota [Rot64].
@@ 2.2.2 Partition algebra
To each $\pi\in\mathrm{Part}\left([2k]\right)$ we associate a diagram consisting of two rows of $k$ vertices labeled from $1,\dots k$ on the top row and from $k+1,\dots,2k$ on the bottom row. We draw an edge between any two vertices $i$ and $j$ if and only if $i\sim j$ in $\pi.$ Any diagram consisting of two rows of $k$ vertices with all connected components complete clearly corresponds to an element $\pi\in\mathrm{Part}\left([2k]\right).$ The product $\pi_{1}\pi_{2}$ is obtained by merging the bottom row of vertices of $\pi_{1}$ with the top row of vertices of $\pi_{2}$ to obtain a new diagram $\tilde{\pi}_{3}$ with three rows of $k$ vertices; adding edges between any vertices that are in the same connected component (where there is not already an edge); removing the middle row of vertices and any adjacent edges to obtain $\pi_{3}\in\mathrm{Part}\left([2k]\right).$ Then we have
@@
Martin and Saleur [MS93] describe some important properties of $P_{k}(n).$
@@
The partition algebra $P_{k}(n)$ is semisimple for any $n\in\mathbb{C},$ unless $n$ is an integer in the range [0,2k1].[0,2k-1].
@@ 2.3 Schur–Weyl–Jones duality
There is an analogue of classical Schur–Weyl duality due to Jones [Jon94], which we refer to as Schur–Weyl–Jones duality.
@@
$\big\langle\left(e_{i_{i}}\otimes\dots\otimes e_{i_{k}}\right)\pi,\ e_{i_{k+1}}\otimes\dots\otimes e_{i_{2k}}\big\rangle=\begin{cases}1&\mathrm{if}\ j\sim l\mathrm{\ in}\ \pi\iff i_{j}=i_{l}\\ 0&\mathrm{otherwise.}\end{cases}$
@@
The symmetric group $S_{n}$ acts on $\left(\mathbb{C}^{n}\right)^{\otimes k}$ on the left via the diagonal permutation action, i.e. for $g\in S_{n}$ and basis vector $e_{i_{i}}\otimes\dots\otimes e_{i_{k}}\in\left(\mathbb{C}^{n}\right)^{\otimes k},$
@@ 2.4 Refinement
We now present a refinement of Schur–Weyl–Jones duality due to Sam and Snowden [SS15, Section 6.1.3]. We will assume that $k\in\mathbb{Z}_{>0}$ and $n\geq 2k$.
@@
$\overset{\mathrm{def}}{=}$ $e_{i_{1}}\otimes\dots\otimes e_{i_{j-1}}\otimes e_{i_{j+1}}\otimes\dots\otimes e_{i_{k}}.$
@@
the complex linear span of all kk–tuples of distinct elements.
@@
$A_{k}(n)=D_{k}(n)\cap\bigcap_{j=1}^{k}\ker\left(T_{j}\right).$
@@
$A_{k}(n)=\bigcap_{\pi\in\ker\left(R\right)}\ker\left(\pi\right).$
@@
This extends linearly to define an action of $\mathbb{C}\left[S_{k}\right]$ on $\left(\mathbb{C}^{n}\right)^{\otimes k}$ (and suitable subspaces). We then view $A_{k}(n)$ as a representation of $S_{n}\times S_{k},$ where once again $g\in S_{n}$ acts via the diagonal permutation action and $\sigma\in S_{k}$ acts via $p_{\iota\left(\sigma^{-1}\right)}.$ We denote this representation
@@
The version of Schur–Weyl duality on $A_{k}(n)$ is the following.
@@
The main result of the author in [Cas25] is a formula for the orthogonal projection from $\left(\mathbb{C}^{n}\right)^{\otimes k}$ to subspaces isomorphic to the irreducible ‘blocks’ in the decomposition of $A_{k}(n).$ Define
@@
$\mathcal{U}_{\lambda^{+}(n)}\overset{\mathrm{def}}{=}\big\langle\Delta\left(g,\sigma\right)\left(\xi_{\lambda}^{\mathrm{norm}}\right):\ g\in S_{n},\ \sigma\in S_{k}\big\rangle_{\mathbb{C}}\cong V^{\lambda^{+}(n)}\otimes V^{\lambda}.$
This leads to the main theorem and a corollary of the author in [Cas25] which will be key for the combinatorial integration method used in this paper. We will write
$\sum_{\pi}^{\leq S_{k}}$
to indicate that we are summing over set partitions $\pi$ that are $\leq\iota(\tau)$ for some $\tau\in S_{k}.$
@@
$\mathcal{Q}_{\lambda,n}:\left(\mathbb{C}^{n}\right)^{\otimes k}\to\mathcal{U}_{\lambda^{+}(n)}$
@@
$\sum_{\pi\in\mathrm{Part}\left([2k]\right)}^{\leq S_{k}}c(n,k,\lambda,\pi)p_{\pi},$
@@
$c(n,k,\lambda,\pi)=\frac{d_{\lambda^{+}(n)}(-1)^{|\pi|+k}}{(n)_{|\pi|}}\sum_{\overset{\tau\in S_{k}}{{\scriptstyle{\scriptscriptstyle\iota(\tau)\geq\pi}}}}\chi^{\lambda}(\tau).$ (5)
@@
$\frac{1}{d_{\lambda}}\mathrm{tr}_{\mathcal{U}_{\lambda^{+}(n)}}\left(g\right)$
$=$ $\frac{1}{d_{\lambda}}\mathrm{tr}_{\left(\mathbb{C}^{n}\right)^{\otimes k}}\left(g,\mathcal{Q}_{\lambda,n}\right).$
If $\pi\leq\iota(\tau)$ for some $\tau\in S_{k},$ then the diagram corresponding to $\pi$ is obtained from that of $\iota(\tau)$ by deleting edges. Let $\mathrm{del}\left(\pi\right)$ be the number of edges deleted (i.e. $\mathrm{del}\left(\pi\right)=|\pi|-k$, so that if $\pi\leq\iota\left(\tau_{1}\right)$ and $\pi\leq\iota\left(\tau_{2}\right),$ then $\mathrm{del}(\pi)$ is independent of whether we count the edges deleted from $\tau_{1}$ or $\tau_{2}$).
@@
This is immediate from the observation that $d_{\lambda^{+}(n)}=O\left(n^{k}\right)$. ∎
@@ 2.5 The Weingarten calculus
An important component of our proof of Theorem 2.10 is the Weingarten calculus. Originally developed by Weingarten [Wei78] and Xu [Xu97] and then more rigorously by Collins [Col03] and Collins and Sniady [CS06], the Weingarten calculus is a method for computing integrals of matrix coefficients using Schur–Weyl duality analogues. Below is the main formulation for the Weingarten calculus for the symmetric group, we direct the reader to the survey [CMN21, Section 3] or indeed any of the aforementioned papers for an in depth discussion. We will denote the matrix coefficients for $g\in S_{n}$ acting diagonally on $\left(\mathbb{C}^{n}\right)^{\otimes k}$ by $g_{ij}.$
@@
For any multi–indices $I=\left(i_{1},\dots,i_{m}\right),$ $J=\left(j_{1},\dots,j_{m}\right)\in[n]^{m},$ with respect to the uniform measure on $S_{n}$, we have
@@ 2.6 $C^{*}$–algebras of free groups
We refer to [NS06] for a thorough introduction. To the free group $F_{r}$ we associate its group algebra $\mathbb{C}\left[F_{r}\right],$ the collection of elements of the form
@@
where each $x(w)\in\mathbb{C}$ and only finitely many are non–zero. One defines a norm on $\mathbb{C}\left[F_{r}\right]$ by
@@
where $\|-\|_{\mathrm{op}}$ is the operator norm on $\ell^{2}\left(F_{r}\right)$. The completion of $\mathbb{C}\left[F_{r}\right]$ in this norm is the $C^{*}$–algebra of $F_{r},$ denoted $C^{*}\left(F_{r}\right)$. The left regular representation of $F_{r}$ is denoted $\lambda$ and admits the reduced norm on $\mathbb{C}\left[F_{r}\right]$ by
@@
One can define a trace $\tau$ on $\mathbb{C}\left[F_{r}\right]$ to be the map
@@ 2.7 Polynomials and random walks
We collect some results and definitions on polynomials and random walks that are required for the analysis in §3. The proofs are readily available in [MdlS24]. See also [CGVTvH24] for more background.
Lemma 2.16 ([MdlS24, Lemma 4.2]).
@@ 2.7.1 Random walks and free groups
We call a symmetric (i.e. $\mu(g)=\mu\left(g^{-1}\right))$ probability measure $\mu$ on the free group $F_{r}$ reasonable if its support is finite, contains the identity element and generates $F_{r}.$
@@
The spectral radius $\rho=\rho(\mu)=\left\|\mu\right\|_{C_{\lambda}^{*}\left(F_{r}\right)}$ measures how fast the probability that a random walk returns to where it started decays. Recall that a proper power in the free group is defined to be an element of the form $u^{d}$ for $u\in F_{r}$ and $d\geq 2.$
Proposition 2.18 ([MdlS24, Proposition 6.1]).
@@
$\mathbb{P}\left(g_{n}\ \mathrm{is\ a\ proper\ power}\right)\leq Cn^{5}\rho^{n}.$
We denote the subspace of elements of $\mathbb{C}\left[F_{r}\right]$ supported in the ball of radius $q$ by $\mathbb{C}_{\leq q}\left[F_{r}\right].$